# ChartMaster: Advancing Chart-to-Code Generation with Real-World Charts and Chart Similarity Reinforcement Learning

### v1

@@ Abstract
The chart-to-code generation task requires MLLMs to convert chart images into executable code. This task faces two major challenges: limited data diversity and insufficient maintenance of visual consistency between generated and original charts during training. Existing datasets mainly rely on seed data to prompt GPT models for code generation, resulting in homogeneous samples. To address this, we propose ReChartPrompt, which leverages real-world, human-designed charts from arXiv papers as prompts instead of synthetic seeds. Using the diverse styles and rich content of arXiv charts, we construct ReChartPrompt-240K, a large-scale and highly diverse dataset. Another challenge is that although SFT effectively improve code understanding, it often fails to ensure that generated charts are visually consistent with the originals. To address this, we propose ChartSimRL, a GRPO-based reinforcement learning algorithm guided by a novel chart similarity reward. This reward consists of attribute similarity, which measures the overlap of chart attributes such as layout and color between the generated and original charts, and visual similarity, which assesses similarity in texture and other overall visual features using convolutional neural networks. Unlike traditional text-based rewards such as accuracy or format rewards, our reward considers the multimodal nature of the chart-to-code task and effectively enhances the model’s ability to accurately reproduce charts. By integrating ReChartPrompt and ChartSimRL, we develop the ChartMaster model, which achieves state-of-the-art results among 7B-parameter models and even rivals GPT-4o on various chart-to-code generation benchmarks. All resources are available at https://github.com/WentaoTan/ChartMaster.
Figure 1: Fine-tuning Qwen2.5-VL-7B with ReChartPrompt data significantly enhances performance, and applying ChartSimRL further elevates it to match or exceed GPT-4o on certain metrics. For better representation, the “Rating” metric in the Plot2Code benchmark is multiplied by 10.
@@ 1 Introduction
The chart-to-code generation task aims to automatically convert chart images into executable code yang2024chartmimic, enabling applications including automated data analysis, report generation, and intelligent question answering zhao2025chartcoder. This task is challenging as it requires accurate visual understanding, cross-modal reasoning, and advanced code synthesis. Although recent advances in Multimodal Large Language Models (MLLMs) show promising results in various vision-language tasks, their performance on chart-to-code generation remains limited due to the unique complexity of charts and the need for precise code output.
@@
Prior work such as ChartCoder zhao2025chartcoder advanced the field by building the large Chart2Code-160K dataset. This dataset is synthesized by guiding GPT-4o hurst2024gpt with predefined chart attributes like chart type, color, and text. While this approach reduces reliance on costly manual annotations and achieves strong performance, relying on predefined attribute seeds can introduce homogeneity and limit variability in the resulting dataset (see Appendix Fig. 5), potentially restricting model generalization to diverse real-world charts.
@@
To address this limitation, we introduce Real-world Chart Prompt Code Generation (ReChartPrompt), a novel automated pipeline that extracts real-world chart images from arXiv papers and uses the Qwen2.5-VL-72B model bai2025qwen2 to generate code. Specifically, we collect 30,071 papers and use their charts as prompts to build a large-scale ReChartPrompt-240K dataset. As the charts in each paper are manually designed by their authors and cover a wide range of research fields, the dataset features diverse expertise and design styles. This rich diversity enables effective generalization to real-world scenarios. As shown in Fig. 1, fine-tuning with ReChartPrompt-240K improves performance on the ChartMimic benchmark by about 30 to 40 points.
@@
While supervised fine-tuning with diverse data can help models generate better chart code, such token-level optimization cannot guarantee that the output charts are visually identical to the references. As shown in Fig. 4, SFT model produces charts closer to the ground truth than the baseline, but differences remain. To address this, we propose ChartSimRL, a reinforcement learning algorithm based on Group Relative Policy Optimization (GRPO) shao2024deepseekmath with a novel chart similarity reward. Specifically, the reward jointly considers attribute similarity—which evaluates the consistency of chart elements such as textual content, numerical values, layout, and color—and visual similarity, which assesses overall visual resemblance using convolutional neural networks (e.g., ResNet he2016deep) to extract and compare visual features. To the best of our knowledge, this is the first reward system designed to effectively capture the multimodal visual-semantic consistency that is crucial for chart-to-code generation. This enables models to produce charts that are not only code-accurate but also faithful to the originals, addressing a critical gap in prior research and supporting more robust generalization to real-world chart reproduction scenarios. As shown in Fig. 1, applying ChartSimRL to SFT-trained models brings performance close to that of GPT-4o.
@@
In summary, we present ChartMaster, an efficient chart-to-code generation framework integrating the ReChartPrompt data generation pipeline and ChartSimRL reinforcement learning strategy. Built on the Qwen2.5-VL-7B model, ChartMaster achieves near GPT-4o performance on multiple benchmarks, advancing automation in chart reproduction. Our main contributions are: (1) ReChartPrompt, an automated data generation method leveraging real academic charts for diverse datasets; (2) ChartSimRL, a reinforcement learning algorithm combining visual and attribute similarity rewards; and (3) ChartMaster-7B, a compact model achieving near GPT-4o performance with only 7B parameters.
@@ 2.1 Multimodal Code Generation
Recently, MLLMs have attracted increasing attention for their code generation capabilities zhang2024humaneval. For example, MMCode li2024mmcode focuses on algorithmic problems in visually rich contexts, where each task is accompanied by one or more images.
Among various code generation tasks, chart-to-code generation has emerged as a critical research area zhao2025chartcoder; yang2024matplotagent. Relevant benchmarks include Design2Code si2024design2code, which evaluates HTML webpage generation using CLIP scores and HTML block comparison metrics, and Plot2Code wu2024plot2code, which assesses both code correctness and visual similarity for comprehensive evaluation. However, since the datasets for Design2Code and Plot2Code are sourced from the web, there is a risk of data leakage, which may compromise the reliability of model evaluation. To address this issue, ChartMimic yang2024chartmimic provides a manually curated dataset of 4,800 chart-code pairs, along with additional fine-grained evaluation metrics.
Despite these benchmarks, large-scale chart-to-code training datasets remain scarce. ChartCoder zhao2025chartcoder addresses this by creating Chart2Code-160K, the first large-scale training set generated by guiding GPT-4o with predefined chart attributes such as type, color, values, and titles. It also uses the “Snippet of Thought” method zheng2023outline; luo2024python to break down code generation into finer steps, significantly enhancing the model’s chart reasoning ability. Nevertheless, the limited diversity of predefined attributes results in relatively homogeneous charts. To overcome this limitation, we propose ReChartPrompt, which leverages real charts from arXiv papers instead of synthetic seeds to generate more diverse and representative chart-code pairs.
@@ 2.2 Reinforcement Learning for MLLMs
Reinforcement learning (RL) has been proven effective in enhancing model capabilities wang2024comprehensive; milani2024explainable. For example, reinforcement learning from human feedback (RLHF) bai2022training and direct preference optimization (DPO) rafailov2023direct align model outputs with human preferences, significantly improving complex reasoning ability and output quality.
Building on these advances, Group Relative Policy Optimization (GRPO) shao2024deepseekmath was developed as a novel RL algorithm that updates policies based on relative rewards computed from groups of samples. For instance, DeepSeek-R1 guo2025deepseek uses simple but effective reward functions based on output accuracy and response format, enabling stable training and fostering emergent reasoning skills such as reflection and “a-ha” moments.
Inspired by DeepSeek-R1’s success, recent work has extended GRPO-based RL to MLLMs tan2025reason; zhang2025r1; peng2025lmm; shen2025vlm. These efforts can be categorized into two groups. The first adapts R1’s approach to MLLMs; for example, Vision-R1 huang2025vision constructs supervised data with reflection for cold-start training, then applies GRPO with accuracy- and format-based rewards. Similarly, MM-EUREKA meng2025mm improves reward design and loss functions, successfully reproducing the visual “aha moment” where the model revisits images and remarks “upon closer inspection.” While impactful, these works mainly target mathematical reasoning tasks. The second category applies GRPO to a wider range of tasks, including visual perception yu2025perception, segmentation liu2025seg, and grounding zhang2025improving, demonstrating GRPO’s robustness and generalizability across diverse domains.
However, to our knowledge, no prior work has applied GRPO to chart-to-code generation. The core difficulty lies in designing an effective reward function that encourages generated code to faithfully reproduce the given chart, both semantically and visually. To this end, we design a novel chart similarity reward that measures the alignment between generated and reference charts, leading to significant improvements in chart reproduction quality.
@@ 3 Method
Figure 2: The overall framework of ChartMaster. (a) Real-world charts from arXiv are curated to create the ReChartPrompt-240K dataset for SFT (SFT is omitted in the figure). (b) The model is further optimized with ChartSimRL. (c) The definition of Chart Similarity; $\mathcal{G}$ denotes the semantic attribute extraction tool; $\mathcal{F}$ is the CNN-based feature extractor; and $\mathbf{f}$ is the extracted feature vector.
@@ 3.1 Using Real-World Charts to Generate Dataset
To improve dataset diversity, we use real-world chart images as input to guide code generation. This approach captures richer styles and content that predefined attribute seeds cannot represent.
@@
(3) Generating code with ReChartPrompt. The Qwen2.5-VL-72B model has demonstrated strong chart-to-code generation capabilities. As an open-source model, it is easily deployed via the vLLM framework kwon2023efficient, making it well-suited for large-scale data generation. We design a set of 20 chart-to-code prompts to enrich instruction diversity, collectively referred to as $\mathbf{P}_{\text{rechart}}$ (see Fig. 7 in Appendix). Below is an example: <<Real-World Chart>>Please generate Python matplotlib code to recreate the picture shown.
@@
Summary. We download 30,071 papers from arXiv, extracting 288,992 chart images. The Qwen2.5-VL-72B model generates code for these charts, and after executing and filtering out failed cases, we obtain 242,479 high-quality triplets, which form the ReChartPrompt-240K dataset. Formally, the dataset is represented as $\mathcal{D}=\{(I_{i},T_{i},Y_{i})\}_{i=1}^{N}$, where $I_{i}$ denotes the chart image, $T_{i}\in\mathbf{P}_{\text{rechart}}$ is the instruction prompt, and $Y_{i}$ is the executable code.
@@ 3.2 Training ChartMaster: SFT and ChartSimRL
Reinforcement Learning with ChartSimRL. While SFT strengthens the model’s basic capability, discrepancies may still exist between the generated charts and the originals (see Fig. 4). To further improve reproduction fidelity, we continue training the model using ChartSimRL. Specifically, for each training sample $(I,T,Y)$, the model samples a group of $M$ candidate codes:
@@
Chart Similarity Reward. Traditional reward functions, such as the accuracy reward used in guo2025deepseek; huang2025vision, primarily assess the consistency between generated text and ground-truth text. However, the chart-to-code task is inherently multimodal, involving both code and generated charts, requiring evaluation of not only semantic correctness but also visual alignment. To this end, we design a novel chart similarity reward as:
@@
Here, $R_{i}^{\mathrm{attr}}$ measures the semantic consistency, and $R_{i}^{\mathrm{vis}}$ captures visual similarity.
@@
Visual Similarity: We use a pretrained ResNet-18 network he2016deep $\mathcal{F}=\{\mathcal{F}_{1},\mathcal{F}_{2},\mathcal{F}_{3},\mathcal{F}_{4}\}$ to extract feature maps from both $I$ and $\hat{I}_{i}$. Here, $\mathcal{F}_{k}(\cdot)\in\mathbb{R}^{C_{k}\times H_{k}\times W_{k}}$ denotes the output feature map of the $k$-th residual block. We extract the feature map and flatten them into vectors like:
@@
Following the GRPO framework shao2024deepseekmath, we update the model by maximizing the clipped surrogate objective with a KL penalty to stabilize training:
@@ 4 Experiment
4.1 Implementation Details
During the collection of arXiv papers, we explicitly exclude papers that are used as benchmarks to avoid potential data leakage. In the data generation stage, we apply a greedy sampling strategy to filter chart data, retaining only images of 12 predefined chart types and discarding all others. Then, we randomly select an instruction from $\mathbf{P}_{\text{rechart}}$ to prompt the Qwen2.5-VL-72B bai2025qwen2 model to generate code via nucleus sampling, with a temperature of 0.1 and a top-p of 0.9.
For training, we use the Qwen2.5-VL-7B model bai2025qwen2 in two stages. In Stage 1, we perform SFT on the entire ReChartPrompt-240K dataset with a learning rate of $2\times 10^{-5}$, batch size 128, and a cosine annealing scheduler for one epoch; the resulting model is saved for Stage 2. In Stage 2, ChartSimRL training is conducted on 10% of the ReChartPrompt-240K dataset, using a smaller learning rate of $5\times 10^{-6}$ and generating $M=4$ candidate codes per sample. Candidate sampling uses temperature 1.0, top-p 1.0, and top-k 80 to encourage diversity. The batch size remains 128 (32 samples $\times$ 4 candidates each).
4.2 Evalution Benchmarks
Following zhao2025chartcoder, we assess the model’s chart-to-code generation performance on multiple benchmarks. ChartMimic Direct Mimic Task yang2024chartmimic: This benchmark includes 600 chart images. GPT-4o scores (0–100) serve as high-level similarity metrics. Additionally, low-level F1 scores for text, layout, chart type, and color are computed from code execution for fine-grained analysis. Plot2Code Direct Asking wu2024plot2code: Metrics include code pass rate, text match rate, and a 10-point GPT-4V visual similarity score, jointly assessing code correctness and visual fidelity. ChartX Chart Redrawing Task xia2024chartx: This benchmark uses GPT-4 (0–5 scale) to evaluate code-generated chart redrawings.
@@ 4.2 Evalution Benchmarks
GeminiProVision team2023gemini 68.2 53.8 53.3 68.2 53.6 3.69 -
Claude-3-opus anthropic2024claude 83.3 60.5 60.1 84.1 57.5 3.80 -
GPT-4V hurst2024gpt 91.2 76.4 78.9 84.1 57.7 5.58 2.63
GPT-4o hurst2024gpt 93.2 79.0 83.5 88.6 56.3 5.71 2.36
@@
ChartAssisstant-13B meng2024chartassisstant - - - - - - 0.82
ChartVLM-L-14B xia2024chartx 19.5 15.8 13.9 - - - 1.58
DeepSeek-VL-7B lu2024deepseek 41.3 19.0 17.6 64.4 32.6 2.26 -
TinyChart-3B zhang2024tinychart 42.5 26.3 25.9 43.2 44.6 2.19 1.89
ChartLlama-13B han2023chartllama 57.5 24.8 28.1 58.4 40.3 2.32 0.94
LLaVA-Next-Mistral-7B li2024llava 59.7 20.7 21.3 72.0 38.7 2.87 -
InternVL2-8B chen2024internvl 61.8 34.4 38.9 77.3 37.1 2.78 1.63
Qwen2-VL-7B wang2024qwen2 67.0 32.9 35.0 68.2 33.8 3.10 1.50
MiniCPM-Llama3-V2.5-8B yao2024minicpm 80.3 36.6 42.1 76.3 37.3 2.61 1.66
Qwen2-VL-72B wang2024qwen2 73.3 54.4 50.9 72.0 53.4 4.26 1.69
InternVL2-Llama3-76B chen2024internvl 83.2 54.8 62.2 85.6 46.6 3.89 1.74
Qwen2.5-VL-72B bai2025qwen2 88.5 72.7 71.8 84.8 68.4 6.83 2.52
ChartCoder-7B zhao2025chartcoder 91.4 77.4 74.0 87.9 54.5 4.50 2.09
Qwen2.5-VL-7B (Baseline) bai2025qwen2 65.5 39.9 34.2 67.4 43.8 4.60 2.18
ChartMaster-7B 93.8 78.2 77.3 88.2 62.6 5.65 2.46

### v2

@@ Abstract
The chart-to-code generation task requires MLLMs to convert chart images into executable code. This task faces two main challenges: limited data diversity and the difficulty of maintaining visual consistency between generated charts and the original ones. Existing datasets mainly rely on synthetic seed data to prompt GPT models for code generation, resulting in homogeneous samples that limit model generalization to real-world chart styles. To address this, we propose ReChartPrompt, leveraging real-world, human-designed charts extracted from arXiv papers as prompts. By harnessing the rich content and diverse visual styles of arXiv charts, we construct ReChartPrompt-240K, a large-scale and highly diverse dataset that better reflects realistic chart variations. For the second challenge, although SFT improves code understanding by optimizing next-token prediction, it does not provide direct supervision on visual features. As a result, it often fails to guarantee that the generated charts visually match the original ones. To address this, we propose ChartSimRL, a GRPO-based reinforcement learning algorithm guided by a novel chart similarity reward. This reward consists of two components: attribute similarity, which measures the overlap of chart attributes like layout and color between the generated and original charts, and visual similarity, which evaluates overall visual features, including texture, using convolutional neural networks. Unlike traditional text-based rewards, our reward accounts for the multimodal nature of the chart-to-code generation task, significantly enhancing the model’s ability to accurately reproduce charts. Integrating ReChartPrompt and ChartSimRL, we develop the ChartMaster model, achieving SOTA results among 7B-parameter models and rivaling GPT-4o on various chart-to-code benchmarks. All resources are available at https://github.com/WentaoTan/ChartMaster.
@@ 1 Introduction
The chart-to-code generation task aims to automatically convert chart images into executable code (Yang et al., 2024a), enabling applications including automated data analysis, report generation, and intelligent question answering (Zhao et al., 2025). This task is challenging as it requires accurate visual understanding, cross-modal reasoning, and advanced code synthesis. Although recent advances in Multimodal Large Language Models (MLLMs) show promising results in various vision-language tasks, their performance on chart-to-code generation remains limited due to the unique complexity of charts and the need for precise code output.
@@
Prior work, such as ChartCoder (Zhao et al., 2025), advanced the field by building the large Chart2Code-160K dataset. This dataset is synthesized by guiding GPT-4o (Hurst et al., 2024) with predefined chart attributes like chart type, color, and text. While this approach reduces the need for costly manual annotations and achieves strong performance, relying on predefined attribute seeds can introduce homogeneity and limit variability in the resulting dataset (see Appendix Fig. 5), potentially restricting model generalization to diverse real-world charts.
Figure 1: Performance comparison on three benchmarks. Our method outperforms ChartCoder-7B (Zhao et al., 2025), and matches or exceeds GPT-4o on certain metrics. For better representation, the “Rating” metric in the Plot2Code benchmark is multiplied by 10.
@@
To address this limitation, we introduce Real-world Chart Prompt Code Generation (ReChartPrompt), a novel automated pipeline that extracts real chart images from arXiv papers and leverages the Qwen2.5-VL-72B model (Bai et al., 2025) to generate corresponding code. By collecting 30,071 papers and utilizing their author-designed charts as prompts, we construct ReChartPrompt-240K, a large-scale dataset comprising 240K chart–code pairs. Since these charts originate from papers across diverse research fields and exhibit a wide variety of design styles, the dataset captures rich visual and semantic diversity, as illustrated in Fig. 5. This heterogeneity enables effective generalization to real-world scenarios.
@@
While supervised fine-tuning with diverse data can help models generate better chart code, such next-token prediction alone does not ensure the output charts are visually faithful to the references. As shown in Fig. 4, the SFT model produces charts closer to the ground truth than the baseline, but noticeable discrepancies remain in color, element positioning, and other visual attributes. To address this, we propose ChartSimRL, a reinforcement learning algorithm based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024), guided by a novel chart similarity reward. Specifically, the reward jointly considers (1) attribute similarity that evaluates the consistency of chart elements such as textual content, numerical values, layout and color, and (2) visual similarity, which assesses holistic visual resemblance using convolutional neural networks (e.g., ResNet (He et al., 2016)) to extract and compare visual features. To the best of our knowledge, this is the first reward system that explicitly enforces multimodal visual-semantic consistency for chart-to-code generation. By encouraging models to produce code that renders charts both semantically accurate and visually faithful, we address a critical gap in prior research and support more robust generalization to real-world chart reproduction.
@@
In summary, we introduce ChartMaster, an efficient framework for chart-to-code generation that combines the ReChartPrompt data generation pipeline with the ChartSimRL reinforcement learning strategy. Our key contributions are: (1) ReChartPrompt, an automated method for generating diverse datasets from real-world charts; (2) ChartSimRL, a reinforcement learning algorithm that uses both visual and attribute similarity to improve output; and (3) ChartMaster-7B, a compact model that delivers near GPT-4o performance with only 7 billion parameters. Fig. 1 highlights its efficiency and effectiveness.
@@ 2.1 Multimodal Code Generation
Multimodal large language models (MLLMs) have recently demonstrated strong capabilities in code generation (Zhang et al., 2024a; Tan et al., 2025b). Notably, MMCode (Li et al., 2024b) targets algorithmic problems embedded in visually rich contexts, where tasks are accompanied by one or more images.
Among multimodal code generation tasks, chart-to-code translation has emerged as a critical challenge (Yang et al., 2024b). Existing benchmarks include Design2Code (Si et al., 2024), which evaluates HTML generation using CLIP scores (Radford et al., 2021) and structural HTML metrics, and Plot2Code (Wu et al., 2024), which assesses both code correctness and visual fidelity. However, since the datasets for Design2Code and Plot2Code are sourced from the web, there is a risk of data leakage, which may compromise the reliability of model evaluation. To address this issue, ChartMimic (Yang et al., 2024a) provides a manually curated dataset of 4,800 chart-code pairs, along with additional fine-grained evaluation metrics.
Despite these benchmarks, large-scale chart-to-code training datasets remain scarce. ChartCoder (Zhao et al., 2025) addresses this by creating Chart2Code-160K, the first large-scale training set generated by guiding GPT-4o with predefined chart attributes such as type, color, values, and titles. It further employs the “Snippet of Thought” strategy (Zheng et al., 2023; Luo et al., 2024) to decompose code generation into structured steps, significantly boosting chart reasoning. Yet, reliance on fixed attributes limits chart diversity. In contrast, our ReChartPrompt leverages real-world charts from arXiv papers as prompts, yielding more diverse and representative chart–code pairs.
@@ 2.2 Reinforcement Learning for MLLMs
Reinforcement learning (RL) effectively enhances model capabilities Wang et al. (2024b); Milani et al. (2024); Tan et al. (2025c). For example, RL from human feedback (RLHF) Bai et al. (2022) and direct preference optimization (DPO) Rafailov et al. (2023) aligned model outputs with human preferences, improving complex reasoning and output quality. Building on these advances, Group Relative Policy Optimization (GRPO) Shao et al. (2024) was proposed as a novel RL algorithm that updated policies using relative rewards computed from groups of samples. DeepSeek-R1 Guo et al. (2025) employed simple yet effective rewards based on output accuracy and response format, which enabled stable training and emergent reasoning such as reflection and “a-ha” moments.
Inspired by DeepSeek-R1’s success, recent work extended GRPO-based RL to MLLMs (Tan et al., 2025a; Zhang et al., 2025b; Peng et al., 2025; Shen et al., 2025) in two main directions. The first adapts R1’s method to MLLMs—for instance, Vision-R1 (Huang et al., 2025) uses SFT data with reflection for cold-start training and applies GRPO with accuracy- and format-based rewards. Similarly, MM-EUREKA (Meng et al., 2025) refines reward design and loss functions, successfully reproducing the visual “aha moment,” where the model revisits images “upon closer inspection.” These works primarily focus on mathematical reasoning tasks. The second direction applies GRPO to broader tasks such as visual perception (Yu et al., 2025), segmentation (Liu et al., 2025), and grounding (Zhang et al., 2025a), demonstrating its robustness and generalizability across domains.
However, to our knowledge, GRPO has not been applied to chart-to-code generation, mainly due to the challenge of designing reward functions that encourage generated code to faithfully reproduce charts both semantically and visually. We address this by proposing a novel chart similarity reward, significantly improving chart reproduction quality.
@@ 3 Method
Figure 2: The overall framework of ChartMaster. (a) Real-world charts from arXiv are curated to create the ReChartPrompt-240K dataset for SFT (SFT is omitted in the figure). (b) The model is further optimized with ChartSimRL. (c) The definition of Chart Similarity: $\mathcal{G}$ denotes the semantic attribute extraction tool; $\mathcal{F}$ is the CNN-based feature extractor; and $\mathbf{f}$ is the extracted feature vector.
@@ 3.1 Using Real-World Charts to Generate Dataset
To improve dataset diversity, we use real-world chart images as input to guide code generation, as shown in Fig. 2 (a). This approach captures richer styles and content that predefined attribute seeds cannot represent.
@@
(3) Generating code with ReChartPrompt. The Qwen2.5-VL-72B model has demonstrated strong chart-to-code generation capabilities. As an open-source model, it is easily deployed via the vLLM framework (Kwon et al., 2023), making it well-suited for large-scale data generation. We design a set of 20 chart-to-code prompts to enrich instruction diversity, collectively referred to as $\mathbf{P}_{\text{rechart}}$ (see Fig. 7 in Appendix). Below is an example: <<Real-World Chart>>Please generate Python matplotlib code to recreate the picture shown.
@@
Summary. We download 30,071 papers from arXiv and extract their figures, filtering out non-chart ones to obtain 288,992 chart images. Using these charts, the Qwen2.5-VL-72B model generates corresponding code. After executing the generated code and removing failed cases, we collect 242,479 high-quality triplets that constitute the ReChartPrompt-240K dataset. Formally, the dataset is defined as $\mathcal{D}=\{(I_{i},T_{i},Y_{i})\}_{i=1}^{N}$, where $I_{i}$ represents a chart image, $T_{i}\in\mathbf{P}_{\text{rechart}}$ is the instruction prompt, and $Y_{i}$ denotes the executable code. Notably, all real-world chart data and generation models employed in this process are open-source, ensuring minimal cost and excellent scalability.
@@ 3.2 Training ChartMaster: SFT and ChartSimRL
Reinforcement Learning with ChartSimRL. While SFT strengthens the model’s basic capability, discrepancies may still exist between the generated charts and the originals (see Fig. 4). To further improve reproduction fidelity, we continue training the model using ChartSimRL, as illustrated in Fig. 2 (b). Specifically, for each training sample $(I,T,Y)$, the model samples a group of $M$ candidate codes:
@@
Chart Similarity Reward. Traditional reward functions, such as the accuracy reward used in (Guo et al., 2025; Huang et al., 2025), primarily assess the consistency between generated text and ground-truth text. However, the chart-to-code task is inherently multimodal, involving both code and generated charts, requiring evaluation of not only semantic correctness but also visual alignment. To this end, we design a novel chart similarity reward as:
@@
Here, $R_{i}^{\mathrm{attr}}$ measures the semantic consistency, and $R_{i}^{\mathrm{vis}}$ captures visual similarity (see Fig. 2 (c)).
@@
Visual Similarity: We use a pretrained ResNet-18 network (He et al., 2016) $\mathcal{F}=\{\mathcal{F}_{1},\mathcal{F}_{2},\mathcal{F}_{3},\mathcal{F}_{4}\}$ to extract feature maps from both $I$ and $\hat{I}_{i}$. Here, $\mathcal{F}_{k}(\cdot)\in\mathbb{R}^{C_{k}\times H_{k}\times W_{k}}$ denotes the output feature map of the $k$-th residual block. We extract the feature map and flatten them into vectors like:
@@
Following the GRPO framework (Shao et al., 2024), we update the model by maximizing the clipped surrogate objective with a KL penalty to stabilize training:
@@
Summary. ReChartPrompt and ChartSimRL have been effectively integrated into the ChartMaster framework. This framework not only leverages real-world data for enhanced data diversity but also employs a novel algorithm to ensure visual and semantic alignment in chart reproduction. Consequently, ChartMaster stands as a comprehensive solution for the chart-to-code generation task, demonstrating marked improvements in performance and generalization capabilities.
@@ 4 Experiment
GeminiProVision (Team et al., 2023) 68.2 53.8 53.3 68.2 53.6 3.69 -
Claude-3-opus (Anthropic, 2024) 83.3 60.5 60.1 84.1 57.5 3.80 -
GPT-4V (Hurst et al., 2024) 91.2 76.4 78.9 84.1 57.7 5.58 2.63
GPT-4o (Hurst et al., 2024) 93.2 79.0 83.5 88.6 56.3 5.71 2.36
@@
ChartAssisstant-13B (Meng et al., 2024) - - - - - - 0.82
ChartVLM-L-14B (Xia et al., 2024) 19.5 15.8 13.9 - - - 1.58
DeepSeek-VL-7B (Lu et al., 2024) 41.3 19.0 17.6 64.4 32.6 2.26 -
TinyChart-3B (Zhang et al., 2024b) 42.5 26.3 25.9 43.2 44.6 2.19 1.89
ChartLlama-13B (Han et al., 2023) 57.5 24.8 28.1 58.4 40.3 2.32 0.94
LLaVA-Next-Mistral-7B (Li et al., 2024a) 59.7 20.7 21.3 72.0 38.7 2.87 -
InternVL2-8B (Chen et al., 2024) 61.8 34.4 38.9 77.3 37.1 2.78 1.63
Qwen2-VL-7B (Wang et al., 2024a) 67.0 32.9 35.0 68.2 33.8 3.10 1.50
MiniCPM-Llama3-V2.5-8B (Yao et al., 2024) 80.3 36.6 42.1 76.3 37.3 2.61 1.66
Qwen2-VL-72B (Wang et al., 2024a) 73.3 54.4 50.9 72.0 53.4 4.26 1.69
InternVL2-Llama3-76B (Chen et al., 2024) 83.2 54.8 62.2 85.6 46.6 3.89 1.74
Qwen2.5-VL-72B (Bai et al., 2025) 88.5 72.7 79.1 84.8 68.4 6.83 2.52
ChartCoder-7B (Zhao et al., 2025) 91.4 77.4 74.0 87.9 54.5 4.50 2.09
Qwen2.5-VL-7B (Baseline) (Bai et al., 2025) 65.5 39.9 40.7 67.4 43.8 4.60 2.18
ChartMaster-7B 93.8 78.2 85.1 88.2 62.6 5.65 2.46